a pastry chef created a 50-ounce sugar solution that was 34% sugar from a 20% sugar solution and from a 40% sugar solution. how much of the 20% sugar solution and how much of the 40% sugar solution were used?
Answer by jorel1380(3719)

x+y=50
.40x+.20y=.34(50)=17
40x+20y=1700
20x+20y=1000
20x=700
x=35
y=15
The chef used 35 oz. of 40% sugar and 15 oz. of 20% sugar..